There are 90 good and 10 defective screws in the box. What is the probability that there are no defective screws among 10 randomly taken screws?
1
Expert's answer
2010-07-07T04:55:09-0400

The total amount of screws in the box is: 90+10 = 100.
The number of ways to select 10 screws from 100 is: N = (C100)^10.
The number of ways to select 10 screws from good ones is: m = (C90)^10.
Thus, the probability of taking 10 good screws is: P = m/N = ((C90)^10)/((C100)^10) = (90!*90!)/(80!*100!) ≈ 0.33